# Broad-Line AGN at : The Black Hole Mass Function and a Connection with Little Red Dots

### v1

@@ Abstract
We present a sample of 50 H detected broad-line active galactic nuclei (BLAGN) at redshifts using data from the CEERS and RUBIES surveys. We select these sources directly from JWST/NIRSpec G395M/F290LP spectra. We use a multi-step pre-selection and a Bayesian fitting procedure to ensure a high-quality sample of sources with broad Balmer lines and narrow forbidden lines. We compute rest-frame ultraviolet and optical spectral slopes for these objects, and determine that 10 BLAGN in our sample are also little red dots (LRDs). These LRD BLAGN, when examined in aggregate, show broader H line profiles and a higher fraction of broad-to-narrow component H emission than non-LRD BLAGN. Moreover, we find that of these objects are intrinsically reddened (), independent of the contributions of emission lines to the broadband photometry. We construct the black hole (BH) mass function at after computing robust observational and line detection completeness corrections. This BH mass function shows broad agreement with both recent JWST/NIRSpec and JWST/NIRCam WFSS based BH mass functions, though we extend these earlier results to log (M. The derived BH mass function is consistent with a variety of theoretical models, indicating that the observed abundance of black holes in the early universe is not discrepant with physically-motivated predictions. The BH mass function shape resembles a largely featureless power-law, suggesting that any signature from black-hole seeding has been lost by redshift 5–6. Finally, we compute the BLAGN UV luminosity function and find good agreement with JWST-detected BLAGN samples from recent works, finding that BLAGN hosts constitute 10% of the total observed UV luminosity at all but the brightest luminosities.
@@ 1 Introduction
It is well known that the majority of galaxies host supermassive black holes (SMBHs) that play critical roles in galaxy evolution across cosmic time. These roles are most evident when these SMBHs actively accrete baryonic material and become active galactic nuclei (AGN; e.g. Heckman & Best 2014; Greene et al. 2020). AGN are classically divided into two categories. In type I AGN, the accretion disk of infalling material is unobscured, allowing an observer to directly probe the emission from the rapidly rotating accretion disk. In type II AGN, the accretion disk is obscured by gas and dust, and signatures from the highly kinetic matter near the central SMBH are hidden. As such, type I AGN are distinctly easier to identify due to their clear highly kinetically broadened emission lines, and are thus often referred to as broad-line AGN (BLAGN). While emission line diagnostics such as the BPT diagram (Baldwin et al., 1981; Kewley et al., 2001; Kauffmann et al., 2003) can successfully distinguish both type I and type II AGN from star-forming galaxies in the local universe, these diagnostics are not yet calibrated nor reliable in the high-redshift universe Garg et al. (2022); Maiolino et al. (2023a); Pacucci et al. (2023); Pacucci & Loeb (2024); although some systems also do appear to match local relations e.g., (Maiolino et al., 2023b; Kokorev et al., 2024a). Despite these challenges, observations of these objects are critical to advancing our understanding of galaxy evolution at early cosmic times.
@@
The challenge of detecting these AGN was removed with the launch of the JWST in late 2021. In the following years, due primarily to JWST’s unprecedented near-infrared sensitivity, enormous progress has been made in detecting BLAGN at , resulting in BLAGN detections at and beyond (e.g. Larson et al., 2023a; Harikane et al., 2023; Maiolino et al., 2023a; Kocevski et al., 2023, 2024; Matthee et al., 2024). These studies have successfully detected BLAGN down to in large area surveys such as CEERS (Finkelstein et al., 2017), UNCOVER (Bezanson et al., 2022), JADES (Eisenstein et al., 2023), EIGER (Kashino et al., 2023), and FRESCO (Oesch et al., 2023).
@@ 2.1 Photometry
The analysis in this paper makes use of JWST/NIRCam photometry in the CEERS and PRIMER/UDS fields. The CEERS-field imaging is from the CEERS team’s v1.0 reduction (Finkelstein et al., in prep; these data will soon be available on the CEERS website), and are supplemented by F090W data on the same field from JWST GO#2234, (PI Banados). The imaging in the UDS field comes from the PRIMER survey (PI Dunlop), and is an internal reduction from the PRIMER team (internal version 0.6). Both fields contain the same set of NIRCam filters (F090W, F115W, F150W, F200W, F277W, F356W, F410M and F444W). We also make use of HST/ACS F606W and F814W imaging in both fields (Grogin et al., 2011; Koekemoer et al., 2011). The photometry catalogs also include HST/WFC3 F105W, F125W, F140W and F160W information, but as this is superseded by the NIRCam imaging, these bands are not used in the analysis.
@@ 2.2 Spectroscopy
We analyze data from the Cycle 1 CEERS survey (JWST ERS#1345, PI Finkelstein; Finkelstein et al., 2017, Finkelstein et al. 2024, in prep) and the Cycle 2 RUBIES program (JWST Cycle 2 GO#4233, PIs de Graaf and Brammer; de Graaff et al., 2023; Wang et al., 2024, de Graaff et al. 2024, in prep).
@@
RUBIES consists of six pointings in the CEERS (EGS) field, and 15 pointings in the PRIMER-UDS field; we make use of the 15 pointings available as of 2024 August 9 (all six in CEERS and 9/15 in PRIMER-UDS). All of these pointings have PRISM and G395M configuration data, although not every object observed has data in both configurations. Each configuration and each pointing used three one-shutter nods and NRSIRS2RAPID readout mode with 65 groups per nod (with a single integration per nod) for a total of 2889 seconds of exposure time.
As there are no G140M or F235M RUBIES data, and to best manage the scope of this investigation, we limit our search for BLAGN to the combined samples of G395M data from both programs and supplement with PRISM data for additional measurements when available. Throughout this work, we construct object names as the combination of the observing program and field (CEERS, RUBIES-EGS, and RUBIES-UDS) and the source’s MSA ID, separated by an hyphen. In total, we examine 677 G395M spectra from CEERS and 3879 G395M/ spectra from RUBIES.
@@ 3.1 H Line Detection and Fitting
Initially, we perform a pre-selection to identify candidate lines of interest (at low computational cost). In this pre-selection, we remove the overall continuum shape by smoothing the spectrum using a 151-element wide median filter ( km s-1 in width, such that even the broadest of emission lines will be unaffected). We subtract this smoothed spectrum from the original to remove the continuum shape. We then examine every pixel in a spectrum and select pixels with a signal-to-noise ratio (S/N) greater than five to search for the peaks of potential emission lines. At each of these pixels, we fit a model broad-line with both narrow and broad components centered on that pixel (in a non-continuum subtracted spectrum) using the curve_fit routine from the scipy package for Python (Virtanen et al., 2020). This model consists of the sum of two degenerate Gaussians that share a common central wavelength with independent amplitudes and full-width-half-maxima (FWHMs), and a vertical offset parameter (zeroth order) to account for any underlying (typically very weak or non-detected in G395M) continuum. In all of our models, we forward-model the effects of instrumental broadening by adding the modeled line width and the spectral resolution of the G395M grating (ranging from 400 km s-1 at 3 m [] to 250 km s-1 at 5 m []) in quadrature such that we may fit the underlying kinematic FWHM of measured lines. As such, hereafter all given FWHM values are inherently corrected for the effects of instrumental broadening. In these fits we use flat uninformative priors for all model parameters and require that the FWHM and amplitude of the two model Gaussians be greater than zero, and that the shared line center be within two pixels ( km s-1, depending on the wavelength) of the detected SNR5 pixel.
@@
We next examine the fitted broad-line model and enforce several criteria to select high-fidelity broad emission lines. First, we require the peak of the combined model (the sum of the peaks of the broad and narrow components, and the continuum offset) is at least 80% of the peak observed flux in the fitted region. We then require that the broader of the two fitted Gaussians has a FWHM greater than 500 km s-1. In later steps, we further restrict this value to 700 km s-1 and introduce additional components to model the [N ii] doublet (assuming that the broad-line is H), but we initially leave this cut more conservative so as to avoid removing potential BLAGN objects before conducting more robust Bayesian fitting. Finally, we require that the broader Gaussian have an amplitude that is at least 50% of the median of the noise array in the fitted region. As before, we deliberately leave this cut quite conservative to avoid removing real BLAGN before our Bayesian fitting. These requirements ensure that a candidate pixel may indeed be the peak of a line with a significant broad component that warrants additional study and more sophisticated and computationally expensive fitting. Finally, we require that the pixel be far enough from the end of the spectral wavelength range such that the fitted broad-line’s flux falls to 10% of the peak line flux before the spectrum ends. This requirement helps to reject candidate lines that are simply the result of noisy data near the ends of the spectral wavelength range. After applying the above methodology to all of the CEERS and RUBIES G395M/ spectra, we discover a population of 1114 candidate broad-lines out of a total of 4556 spectra.
@@
We next re-fit these 1114 candidate broad-lines using Markov-Chain-Monte-Carlo (MCMC) fitting using the emcee package for Python (Foreman-Mackey et al., 2013). Here we model the H and [Nii] lines using a five-component model. We model a H line with a narrow Gaussian component and a broad Gaussian component that share a common line center. We also model [N II]6548,6583 as two Gaussians with a fixed internal amplitude ratio of 1:2.94 (Osterbrock & Ferland, 2006), a common line-width with the narrower H component, and line centers that assume the same redshift as the modelled H line. Finally, we include a simple global flux offset parameter to account for any faintly detected continuum emission. We show an example BLAGN object (RUBIES-EGS-15825) that exhibits significant [N II]6548,6583 emission to best illustrate our BLAGN model in Figure 1. For this model, we impose a number of priors such that we require: (1) positive line fluxes for all line components, (2) the narrow component of H (and [Nii]) is narrower than the broad component and narrower than 500 km s-1, (3) the broad component of H has a FWHM 10,000 km s-1, (4) the H line center is within 250 km s-1 of the previously fitted line center, and (5) the continuum offset is within the 10th and 90th percentile of the flux values in the fitted spectral region. We fit the portion of the full object spectrum that lies within 6500 km s-1 of the previously fitted line center. We initialize 32 MCMC “walkers” using the curve_fit best-fit parameters as starting points (after perturbing these values for each walker to ensure that the walkers are independent) and allow them to explore the parameter space for 5000 steps per walker. To minimize the effects of chain “burn-in” we allow emcee to discard the first 2/3 of the chain, and use the last 1667 steps from each walker for our analysis, resulting in 53344 element posterior distributions for each fitted parameter. During this process, we save the full chains for each fitted parameter so that we may easily calculate the distributions of and thus the uncertainties on derived parameters in later steps in our analysis.
@@
We next require that the MCMC fitted broad components of H have km s-1. We choose this value to provide a balance between probing low-mass AGN while not contaminating our sample with non-AGN objects. We later compensate for potentially missed low-mass AGN in our completeness corrections (§6.1). This cut reduces the sample from 814 candidates, to 730 candidates. We next require that the broad components be detected with a flux S/N, reducing the sample from 730 to 152 objects. This drastic reduction in candidate sample size is by-design, as our pre-selection cuts are deliberately weak. Of the 730 candidates, 434 exhibit a broad component with S/N2, 103 exhibit a broad component with 2S/N3, and 41 exhibit a broad component with 3S/N4. It is unlikely that the S/N3 objects show real broad-line components, and likely only sample noise. While a fraction of the 3S/N4 objects may be BLAGN, but we compensate for these objects in our completeness correction (see §6.1). These requirements produce a near-final sample of 152 spectra. Finally, we require that all spectra pass a visual inspection to reject artifacts and contaminants, and that the objects have clear and unambiguous redshifts. For this, we require that the redshift be based on the presence of at least three strong emission lines (with one typically being H and the others being the [O iii]4959,5007 doublet, the [S ii]6718,6733 doublet, and/or the [S iii]9069,9531 doublet) in the grating or prism spectrum (if available) of an object. We find secure redshifts and observe clean visual inspections for 67 of the 152 remaining broad-lines, with most of the removed candidates () clearly being contamination in the 2D spectra from either artifacts or overlapping spectra, with some others () showing only a single apparent emission line in the G395M spectrum that cannot be used to determine a robust multi-line redshift solution. Finally, the last of the rejected spectra feature blended lines (such as the [S ii] doublet) that were fit as broad-lines. While our broad-line search is designed principally to detect H emitters, our methodology is also capable of selecting other broad-lines. In these cases, the amplitude of the modeled [N ii] doublet is simply minimized during the fitting process. As such, of the 67 broad-line spectra with secure redshifts, 51 spectra feature a broad H line. The other 16 spectra that do not contain broad H consist of one He i line, one Pa- line (first identified in Wang et al. 2024), 11 He i lines, and three Pa- lines. While these objects are interesting in their own right, here we limit our study to the sample of 51 broad-line H spectra at .
@@
One object in our sample was observed by both CEERS and RUBIES and was thus detected as both objects CEERS-2782 and RUBIES-EGS-50052 (and is in fact one of the two BLAGN first discovered in CEERS by Kocevski et al. 2023). Interestingly, the CEERS spectrum (CEERS-2782) only exhibits 35% of the H line flux observed in the RUBIES observation (before applying a flux correction, see §3.2). We attribute this difference in observed flux primarily to severe slit-losses, as the CEERS observation slitlet was quite off-center (see Figure 2). Due to the much higher S/N in the RUBIES-EGS-50052 observation, we use this observation for our analysis of this object. As such, we report a final sample of 50 unique broad-line objects.

### v2

@@ Abstract
We present a sample of 62 H detected broad-line active galactic nuclei (BLAGN) at redshifts using data from the CEERS and RUBIES surveys. We select these sources directly from JWST/NIRSpec G395M/F290LP spectra. We use a multi-step pre-selection and Bayesian fitting to ensure a high-quality sample of sources with broad Balmer lines and narrow forbidden lines. We compute rest-frame ultraviolet and optical spectral slopes for these objects, and determine that 21 BLAGN in our sample are also little red dots (LRDs). These LRD BLAGN, when examined in aggregate, show broader H line profiles and a higher fraction of broad-to-narrow component H emission than non-LRD BLAGN. We find that of LRD BLAGN are intrinsically reddened (), independent of contributions from emission lines to the broadband photometry. We construct the black hole (BH) mass function at after computing robust completeness corrections. This BH mass function shows agreement with recent JWST-based BH mass functions, though we extend these earlier results to log (M. The derived BH mass function is consistent with theoretical models, indicating that the observed abundance of BHs in the early universe may not be discrepant with physically-motivated predictions. The BH mass function shape resembles a largely featureless power-law, suggesting that signatures from BH seeding have been lost by redshift 5–6. Finally, we compute the BLAGN UV luminosity function and find agreement with JWST-detected BLAGN samples from recent works, finding that BLAGN hosts constitute 10% of the total observed UV luminosity at all but the brightest luminosities.
@@ 1 Introduction
It is well known that the majority of galaxies host supermassive black holes (SMBHs) that play critical roles in galaxy evolution across cosmic time. These roles are most evident when these SMBHs actively accrete baryonic material and become active galactic nuclei (AGN; e.g. Heckman & Best 2014; Greene et al. 2020). AGN are classically divided into two categories. In type I AGN, the accretion disk of infalling material is unobscured, allowing an observer to directly probe the emission from the rapidly rotating accretion disk. In type II AGN, the accretion disk is obscured by gas and dust, and signatures from the highly kinetic matter near the central SMBH are hidden. As such, type I AGN are distinctly easier to identify due to their clear highly kinetically broadened emission lines, and are thus often referred to as broad-line AGN (BLAGN). While emission line diagnostics such as the BPT diagram (Baldwin et al., 1981; Kewley et al., 2001; Kauffmann et al., 2003) can successfully distinguish both type I and type II AGN from star-forming galaxies in the local universe, these diagnostics are not yet calibrated nor reliable in the high-redshift universe (e.g., Garg et al., 2022; Brinchmann, 2023; Pacucci et al., 2023; Scholtz et al., 2023; Chisholm et al., 2024; Mazzolari et al., 2024; Pacucci & Loeb, 2024); although some systems also do appear to match local relations (e.g., Maiolino et al., 2023a; Kokorev et al., 2024a). Despite these challenges, observations of these objects are critical to advancing our understanding of galaxy evolution at early cosmic times.
@@
The challenge of detecting these AGN was removed with the launch of the JWST in late 2021. In the following years, due primarily to JWST’s unprecedented near-infrared sensitivity, enormous progress has been made in detecting BLAGN at , resulting in BLAGN detections at and beyond (e.g. Larson et al., 2023a; Harikane et al., 2023; Maiolino et al., 2023b; Kocevski et al., 2023, 2024; Matthee et al., 2024). These studies have successfully detected BLAGN down to in large area surveys such as CEERS (Finkelstein et al., 2017), UNCOVER (Bezanson et al., 2022), JADES (Eisenstein et al., 2023), EIGER (Kashino et al., 2023), and FRESCO (Oesch et al., 2023).
@@ 2.1 Photometry
The analysis in this paper makes use of JWST/NIRCam photometry in the CEERS and PRIMER/UDS fields. The CEERS-field imaging is from the CEERS team’s v1.0 reduction (Finkelstein et al., in prep; these data will soon be available on the CEERS website), and are supplemented by F090W data on the same field from JWST GO#2234, (PI Banados). The imaging in the UDS field comes from the PRIMER survey (PI Dunlop), and is an internal reduction from the PRIMER team (internal version 0.6). The PRIMER imaging data were reduced using the PRIMER Enhanced NIRCam Image Processing Library (PENCIL; Magee et al., in preparation, Dunlop et al. in preparation) software. The astrometry of all the reduced images was aligned to GAIA Data Release 3 (Gaia Collaboration et al., 2023) and stacked to the same pixel scale of 003. Both fields contain the same set of NIRCam filters (F090W, F115W, F150W, F200W, F277W, F356W, F410M and F444W). We also make use of HST/ACS F606W and F814W imaging in both fields (Grogin et al., 2011; Koekemoer et al., 2011). The photometry catalogs also include HST/WFC3 F105W, F125W, F140W and F160W information, but as this is superseded by the NIRCam imaging, these bands are not used in the analysis.
@@ 2.2 Spectroscopy
We analyze data from the Cycle 1 CEERS survey (JWST ERS#1345, PI Finkelstein; Finkelstein et al., 2017, 2025) and the Cycle 2 RUBIES program (JWST Cycle 2 GO#4233, PIs de Graaf and Brammer; de Graaff et al., 2023, 2024).
@@
RUBIES consists of six pointings in the CEERS (EGS) field, and 12 pointings in the PRIMER-UDS field. All of these pointings have PRISM and G395M configuration data, although not every object observed has data in both configurations. Each configuration and each pointing used three one-shutter nods and NRSIRS2RAPID readout mode with 65 groups per nod (with a single integration per nod) for a total of 2889 seconds of exposure time.
As there are no G140M or G235M RUBIES data, and to best manage the scope of this investigation, we limit our search for BLAGN to the combined samples of G395M data from both programs and supplement with PRISM data for additional measurements when available.Throughout this work, we construct object names as the combination of the observing program and field (CEERS, RUBIES-EGS, and RUBIES-UDS) and the source’s MSA ID, separated by an hyphen. In total, we examine 677 G395M spectra from CEERS and 4521 G395M spectra from RUBIES.
@@ 3.1 H Line Detection and Fitting
Initially, we perform a pre-selection to identify candidate lines of interest (at low computational cost). In this pre-selection, we remove the overall continuum shape by smoothing the spectrum using a 151-element wide median filter ( km s-1 in width, such that even the broadest of emission lines will be unaffected). We subtract this smoothed spectrum from the original to remove the continuum shape. We then examine every pixel in a spectrum and select pixels with a signal-to-noise ratio (S/N) greater than five to search for the peaks of potential emission lines. At each of these pixels, we fit a model broad-line with both narrow and broad components centered on that pixel (in a non-continuum subtracted spectrum) using the curve_fit routine from the scipy package for Python (Virtanen et al., 2020). This model consists of the sum of two degenerate Gaussians that share a common central wavelength with independent amplitudes and full-width-half-maxima (FWHMs), and a linear model to account for any underlying (typically very weak or non-detected in G395M) continuum. In all of our models, we forward-model the effects of instrumental broadening by assuming a diffraction-limited point source and adding the modeled line width and the spectral resolution of the G395M grating (230 km s-1) in quadrature such that we may fit the underlying kinematic FWHM of measured lines. As such, hereafter all given FWHM values are inherently corrected for the effects of instrumental broadening. In these fits we use flat uninformative priors for all model parameters and require that the FWHM and amplitude of the two model Gaussians be greater than zero, and that the shared line center be within two pixels ( km s-1, depending on the wavelength) of the detected SNR5 pixel.
@@
We next examine the fitted broad-line model and enforce several criteria to select high-fidelity broad emission lines. First, we require the peak of the combined model (the sum of the peaks of the broad and narrow components, and the continuum offset) is at least 80% of the peak observed flux in the fitted region. We then require that the broader of the two fitted Gaussians has a FWHM greater than 500 km s-1. In later steps, we further restrict this value to 700 km s-1 and introduce additional components to model the [N ii] doublet (assuming that the broad-line is H), but we initially leave this cut more conservative so as to avoid removing potential BLAGN objects before conducting more robust Bayesian fitting. Finally, we require that the broader Gaussian have an amplitude that is at least 50% of the median of the noise array in the fitted region. As before, we deliberately leave this cut quite conservative to avoid removing real BLAGN before our Bayesian fitting. These requirements ensure that a candidate pixel may indeed be the peak of a line with a significant broad component that warrants additional study and more sophisticated and computationally expensive fitting. Finally, we require that the pixel be far enough from the end of the spectral wavelength range such that the fitted broad-line’s flux falls to 10% of the peak line flux before the spectrum ends. This requirement helps to reject candidate lines that are simply the result of noisy data near the ends of the spectral wavelength range. After applying the above methodology to all of the CEERS and RUBIES G395M spectra, we discover a population of 1366 candidate broad-lines out of a total of 5198 spectra.
@@
We next re-fit these 1366 candidate broad-lines using Markov-Chain-Monte-Carlo (MCMC) fitting using the emcee package for Python (Foreman-Mackey et al., 2013). Here we model the H and [Nii] lines using a five-component model. We model a H line with a narrow Gaussian component and a broad Gaussian component that share a common line center. We also model [N II]6548,6583 as two Gaussians with a fixed internal amplitude ratio of 1:2.94 (Osterbrock & Ferland, 2006), a common line-width with the narrower H component, and line centers that assume the same redshift as the modeled H line. Finally, we include a simple linear model to account for any faintly detected continuum emission. We show an example BLAGN object (RUBIES-EGS-15825) that exhibits significant [N II]6548,6583 emission to best illustrate our BLAGN model in Figure 1. For this model, we impose a number of priors such that we require: (1) positive line fluxes for all line components, (2) the narrow component of H (and [Nii]) is narrower than the broad component and narrower than 500 km s-1, (3) the broad component of H has a FWHM 10,000 km s-1, (4) the H line center is within 250 km s-1 of the previously fitted line center, and (5) the continuum offset is within the 10th and 90th percentile of the flux values in the fitted spectral region. We fit the portion of the full object spectrum that lies within 6500 km s-1 of the previously fitted line center. We initialize 32 MCMC “walkers” using the curve_fit best-fit parameters as starting points (after perturbing these values for each walker to ensure that the walkers are independent) and allow them to explore the parameter space for 5000 steps per walker. To minimize the effects of chain “burn-in” we allow emcee to discard the first 2/3 of the chain, and use the last 1667 steps from each walker for our analysis, resulting in 53344 element posterior distributions for each fitted parameter. During this process, we save the full chains for each fitted parameter so that we may easily calculate the distributions of and thus the uncertainties on derived parameters in later steps in our analysis.
@@
We next require that the MCMC fitted broad components of H have km s-1. We choose this value to provide a balance between probing low-mass AGN while not contaminating our sample with non-AGN objects. We later compensate for potentially missed low-mass AGN in our completeness corrections (§6.1). This cut reduces the sample from 1366 candidates, to 1142 candidates. We next require that the broad components be detected with a flux S/N, reducing the sample from 1142 to 176 candidates. This drastic reduction in candidate sample size is by-design, as our pre-selection cuts are deliberately weak. Of the 1142 candidates, 685 exhibit a broad component with S/N2, 215 exhibit a broad component with 2S/N3, and 66 exhibit a broad component with 3S/N4. It is unlikely that the S/N3 objects show real broad-line components, and likely only sample noise. While a fraction of the 3S/N4 objects may be BLAGN, but we compensate for these objects in our completeness correction (see §6.1). These requirements produce a sample of 176 spectra.
We next re-fit fit each object spectrum with a modified version of our above model that excludes the broad H line component as a standard of comparison to test for the significance of the broad line. We compute the Bayesian Information Criterion (BIC) for each model fit. We next use the difference in the BIC between the models to determine which model fits best for each candidate line (with the lower BIC model preferred). We require a BIC in favor of the broad-line model over the narrowline model, suggesting that the inclusion of the broad-line significantly improves the overall spectral fit without overfitting or adding unnecessary components to the model. This requirement reduces the sample of 176 objects to 162 objects, implying that the previously fitted broad-line components in the rejected 14 objects were not significant.
Finally, we require that all spectra pass a visual inspection to reject artifacts and contaminants, and to verify the objects have clear and unambiguous redshifts. For this, we require that the redshift be based on the presence of at least three strong emission lines (with one typically being H and the others being the [O iii]4959,5007 doublet, the [S ii]6718,6733 doublet, and/or the [S iii]9069,9531 doublet) in the grating or prism spectrum (if available) of an object. We find secure redshifts and observe clean visual inspections for 79 of the 162 remaining objects, with most of the removed candidates () clearly being contamination in the 2D spectra from either artifacts or overlapping spectra, with some others () showing only a single apparent emission line in the G395M spectrum that cannot be used to determine a robust multi-line redshift solution. Finally, the last of the rejected spectra feature blended lines (such as the [S ii] doublet) that were fit as broad-lines. We stress that we do not evaluate the significance of the the broad-line using this visual inspection, rather we only reject objects that are clearly artifacts or mis-fit doublets.
While our broad-line search is designed principally to detect H emitters, our methodology is also capable of selecting other broad-lines. In these cases, the amplitude of the modeled [N ii] doublet is simply minimized during the fitting process. As such, of the 79 broad-line spectra with secure redshifts, 63 spectra feature a broad H line. The other 16 spectra that do not contain broad H consist of one Pa- line (first identified in Wang et al. 2024a), seven He i lines, three [S iii] lines, and five Pa- lines. While these objects are interesting in their own right, here we limit our study to the sample of 63 broad-line H spectra at .
@@
One object in our sample was observed by both CEERS and RUBIES and was thus detected as both objects CEERS-2782 and RUBIES-EGS-50052 (and is in fact one of the two BLAGN first discovered in CEERS by Kocevski et al. 2023). Interestingly, the CEERS spectrum (CEERS-2782) only exhibits 35% of the H line flux observed in the RUBIES observation (before applying a flux correction, see §3.2). We attribute this difference in observed flux primarily to severe slit-losses, as the CEERS observation slitlet was quite off-center (see Figure 2). Due to the much higher S/N in the RUBIES-EGS-50052 observation, we use this observation for our analysis of this object. As such, we report a final sample of 62 unique broad-line objects.